expand f(z)=z/(z-1)(2-z) in a laurent series valid 1<|z|<2


1
Expert's answer
2021-09-14T11:19:28-0400

"\dfrac{z}{(z-1)(2-z)}=\dfrac{A}{z-1}+\dfrac{B}{2-z}=\dfrac{2A-Az+Bz-B}{(z-1)(2-z)}"

"z^0: 0=2A-B"


"z^1: 1=-A+B"


"A=1, B=2"

"f(z)=\dfrac{z}{(z-1)(2-z)}=\dfrac{1}{z-1}+\dfrac{2}{2-z}"

Geometric series


"\dfrac{1}{z-1}=\dfrac{1}{z}(\dfrac{1}{1-\dfrac{1}{z}})=\displaystyle\sum_{i=1}^n\dfrac{1}{z^n}, |z|>1"

"\dfrac{2}{2-z}=\dfrac{1}{1-\dfrac{z}{2}}=\displaystyle\sum_{i=0}^n\dfrac{z^n}{2^n}, |z|<2"

"\dfrac{z}{(z-1)(2-z)}=1+\displaystyle\sum_{i=1}^n\bigg(\dfrac{1}{z^n}+\dfrac{z^n}{2^n}\bigg),1<|z|<2"